OpenStudy (anonymous):

13. The area of a triangle with base 19 in. and height 11 in. is _____. Be sure to include units. (1 point)

OpenStudy (zale101):

Area = b*h/2

OpenStudy (zale101):

B=base H=hieght

OpenStudy (vjordillo):

Substitute the values in this equation \[\frac{ 1 }{ 2 }(b)(h)\]

OpenStudy (zale101):

now, plug in and solve

OpenStudy (anonymous):

so i do 19 * 11 and then divide by 2? @Zale101

OpenStudy (calculusxy):

Yes. That's right.

OpenStudy (zale101):

the formula says Base X Height divide by 2

OpenStudy (zale101):

what u did was correct, let me know if u got the answer

OpenStudy (anonymous):

i got 104.5 i dont think i did it right.@Zale101

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

Area of a triangle = base * height /2 = 19 in * 11 in = 209 in/2 = 104.5 in^2.

OpenStudy (zale101):

correct :)

OpenStudy (anonymous):

@Zale101 Thanks for working out the problem with and not just giving me the answer. Can you help me with one more please?

OpenStudy (calculusxy):

Not Zale101 though here. But I will try to help :)

OpenStudy (anonymous):

@calculusxy ok cool thanks.14. The area of a circle with a radius of 15 in., to the nearest tenth is _____. Be sure to include units

OpenStudy (calculusxy):

A radius is half of the diameter. The formula for findning the area of a circle is: d*3.14=area. D=diametre

OpenStudy (vjordillo):

Area of a circle is \[A =\pi r ^{2}\] then plug in the value of r

OpenStudy (calculusxy):

So the radius is 15, meaning that u have to multiply 15*15=225. then 225*3.14=706.858347058. If u round that then it will be 706.9in.

OpenStudy (zale101):

thanks @calculusxy :)

OpenStudy (calculusxy):

no problem :)

OpenStudy (anonymous):

oh man thanks for the help guys.! sorry @calculusxy it took me so long to reply i had to leave and run errands.